OpenStudy (anonymous):

You are driving from town A to town B. From town A you go 3 miles West and 4 miles South. What is the distance between the two towns?

OpenStudy (anonymous):

5 miles. 3^2+4^2=5^2

OpenStudy (anonymous):

1) 1 2) 7 3) 5 4) 2

OpenStudy (anonymous):

3)

OpenStudy (anonymous):

5 miles

OpenStudy (anonymous):

3 *3 + 4*4 =25 quare route of 25 is 5